• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z.-E. Meziani

    Log entry time 00:08:50 on July22,2003

    Entry number 105998

    keyword=Swing Shift Summary

    Start of Shift: 4:00PM
    Still in production runs kin5.12

    run numbers
    2648/22648/3517 16:01-16:14 3M
    Change momentum of HRS right and septum to 1.603 GeV/c
    Start to rotate the field and perform an NMR and EPR measurements.
    17:05 Finished NMR and EPR
    17:10 Move target to reference cell
    17:14 empty reference cell run at 10uA. run# 2649 1.8M
    17:40 empty reference cell run at 2 uA. run# 2650 0.36 M
    17:55 Bleedthrough measurement OK 1.85%
    18:00 N2 target at 140psig.
    Start N2 run# 2651
    run# 2652 (too high a CODA rate)
    run# 2653 (good run)
    18:30 Eugene and Sasha performed the Moller measurement
    Polarization 74% corrected for bleedthrough

    21:28 Start production runs in kin5.13

    2654/22654/3523 21:28-21:41 Total 3 M
    2655/22655/3524 21:42-21:57 Total 6 M
    2656/22656/3525 21:58-22:13 Total 9 M
    2657/22657/3526 22:14-22:27 Total 12M
    2658/22658/3526 22:29-22:44 Total 15M
    2659/22659/3527 22:45-22:58 Total 18M
    2660/22660/3527 22:59-23:13 Total 21 M
    2661/22661/3528 23:14-23:27 Total 24M
    2662/22662/3529 23:28-23:34 Total 25 M

    Asked MCC to remove beam half wave plate

    23:40 Start new production in Kin5.13

    2663/22663/3531 23:40-23:53 Total 3M
    2664/22664/3532 23:54-00:08 Total 6M

    00:09 New shift crew.